# Splitting and gluing in sine-dilaton gravity: matter correlators and the wormhole Hilbert space

### v1

@@ 1 Introduction
As a first step towards understanding bulk physics, in [14, 15] the chord diagrams were argued to be the right language for the emergence of the bulk geometry. Indeed, [15] constructed a two-sided chord Hilbert space defined on the two-sided ”wormhole” Cauchy slice, with extensions to include matter operators, and [16] further analyzed the symmetry algebra acting on this Hilbert space. Later, a series of works [17, 18, 19, 20] proposed a concrete holographic dual of DSSYK, which is called the ”sine-dilaton gravity”. The sine-dilaton gravity was shown to reproduce the dynamics of DSSYK, also referred to as q-Schwarzian quantum mechanics [17, 21], and its canonical quantization reproduces the two-sided chord Hilbert space of [16] . Crucially in this construction, the length in the bulk becomes positive and discrete after gauging the shift symmetry of the conjugate momentum [19]. Besides the connection to DSSYK, the sine-dilaton gravity itself has some very unusual features for a gravitational theory, such as bounded spectrum and the resulting good UV properties, and therefore merits investigation beyond its connection to the SYK model.
In this work, we take the next step in exploring this duality, by including bulk matter fields and reproducing general correlation functions of boundary-inserted operators from the gravitational perspective. A nontrivial example is the out-of-time ordered correlator (OTOC). In DSSYK, the OTOC has been calculated using the Wick contractions in the chord diagram [13], where it gives the 6j-symbol of the quantum group $\mathcal{U}_{q}(su(1,1))$. Notably, this 6j-symbol leads to sub-maximal chaos, which is related to the ”fake thermal circle” of DSSYK [16]. On the gravity side, a direct calculation of OTOC is more subtle. For example, previous OTOC calculations in JT gravity rely on boundary-bulk mappings or bootstrap techniques [22, 23, 24, 25]. However, in sine-dilaton gravity, a more straightforward bulk calculation can be expected – since the bulk geometry is discrete, the separation of different bulk regions in the OTOC contour is less ambiguous, as we will see below.
@@ 1.1 Summary
The motivation to investigate this procedure originates from a natural bulk perspective: a boundary-to-boundary particle worldline on the Euclidean $\mathrm{AdS}_{2}$ disk divides the disk to two regions and can be considered as an end-of-the-world (EOW) brane for each region. When the time direction is chosen such that the Cauchy slices are orthogonal to the EOW brane, this brane acts as a defect operator, leading to a twisted Hilbert space, which we denote by $\mathcal{H}_{\mathrm{\Delta}}$, with $\Delta$ being one of the brane parameter corresponding to the scaling dimension of the inserted boundary matter operator. The canonical quantization result of this twisted Hilbert space was established in [20], where the resolution of the identity operator takes the form
$\textbf{Twisted Hilbert space {$\mathcal{H}_{\mathrm{\Delta}}$}}:\quad\mathds{1}=\int_{0}^{\pi}d\theta\leavevmode\nobreak\ \rho_{\Delta,\bar{\alpha}}(\theta)|\theta\rangle_{\Delta,\bar{\alpha}}{{}_{\Delta,\bar{\alpha}}}\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|.$ (1)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}\quad=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}.$ (2)
@@
    $|\Delta;\theta_{L},\theta_{R}\rangle_{0}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=|{\theta_{L}}\rangle\otimes^{\theta_{L}}|{\theta_{R}}\rangle_{\Delta,\theta_{L}},$ (3)
@@
    $|\Delta;\theta_{L},\theta_{R}\rangle_{1}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{R},\bar{\alpha}}$ (5)
@@
$|\Delta_{1},...,\Delta_{n};\theta_{1},...,\theta_{n+1}\rangle=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$},$ (7)
@@
$\mathrm{OTOC}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}.$ (10)
@@
$\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\,\rho(\bar{\alpha}_{1})\,\rho(\bar{\alpha}_{2})\quad\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}=\leavevmode\hbox{$\vbox{\hbox{\resizebox{}{}{{\leavevmode\hbox{\set@color{}}}}}}$}.$ (11)
@@ 1.2 Sine-dilaton gravity setup
$\int\mathcal{D}g\mathcal{D}\Phi\leavevmode\nobreak\ \mathrm{exp}\left(\frac{1}{2|\log q|}\left\{\frac{1}{2}\int d^{2}x\sqrt{g}\left(\Phi R+2\sin(\Phi)\right)+\int_{\mathrm{bdy}}du\sqrt{h}\leavevmode\nobreak\ \Phi K-i\int_{\mathrm{bdy}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}\right\}\right),$ (12)
@@
$\sqrt{h}\leavevmode\nobreak\ e^{i\Phi/2}\Big{|}_{\Phi=\Phi_{\mathrm{bdy}}}=\frac{i}{2|\log q|},\quad\Phi_{\mathrm{bdy}}=\frac{\pi}{2}+i\infty.$ (13)
@@
$L=\oint_{\mathrm{slice}}ds\leavevmode\nobreak\ e^{-i\Phi/2}=\oint_{\mathrm{slice}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}.$ (17)
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\leavevmode\nobreak\ \rho(\theta)|\theta\rangle\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|,$ (21)
@@
$Z_{\mathrm{SD}}(\beta)$ $=\langle n=0|e^{-\beta\hat{H}_{\mathrm{SD}}}|n=0\rangle=\int_{0}^{\pi}d\theta\leavevmode\nobreak\ \rho(\theta)\leavevmode\nobreak\ e^{\beta\frac{\cos(\theta)}{\hbar}},$ (25)
@@ 2 Matter fields in sine-dilaton gravity
$\langle\mathcal{O}(x_{1})\mathcal{O}(x_{2})\rangle=\int\mathcal{D}x\exp\left(-\frac{m}{\hbar}\int_{x(u_{i})=x_{1}}^{x(u_{f})=x_{2}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}\,+\mathrm{c.t.}\right)=\ \exp\left(-\Delta L(x_{1},x_{2})\right),$ (27)
where we emphasize that the action $\frac{m}{\hbar}\int_{x(u)}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}$ is formally divergent as the bulk operator approaches the boundary, and $L(x_{1},x_{2})$ is a suitably renormalized geodesic length between the boundary points $x_{1}$ and $x_{2}$ in $\mathrm{AdS}_{2}$, obtained after adding the appropriate counter term (c.t.) [^2] The precise form of the counter term, as well as the procedure for obtaining the effective action on the geodesic, is currently unknown. The difficulty in identifying it arises from the fact that the boundary of sine-dilaton gravity is essentially non-geometric and singular [31], as $\mathrm{AdS}_{2}$ patch of sine-dilaton gravity is defined along a complex contour that extends to $i\infty$ and the classical curvature diverges. In this paper, we take the conjectured dictionary (28), inspired by the quantum group structure and the JT gravity result [32], where the effective action $e^{-\Delta L(x_{1},x_{2})}$ is in fact the exact result for any $\Delta$, as the definition of this counter term. It would be interesting and important to investigate this point in the future., and $\Delta$ is the scaling dimension of the operator inserted at the boundary. The relation of the scaling dimension to the particle mass is conjectured to be [20, 17]
@@ 2.1 Splitting a particle worldline in $\mathrm{AdS}_{2}$
$-im\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}+\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ \Phi K_{L}+\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ \Phi K_{R}.$ (31)
@@
$\exp\left(-\frac{m}{\hbar}\int_{x(u)}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}\right)\rightarrow\exp\left(-\Delta L\right)=\exp\left(-(\Delta_{L}+\Delta_{R})L\right)$
$\rightarrow\exp\left(-\frac{m_{L}}{\hbar}\int_{x(u)}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}-\frac{m_{R}}{\hbar}\int_{x(u)}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}\right)$ (32)
where in the first line, the arrow indicates the use of holographic renormalization and summing over all possible shell locations to obtain the finite action $e^{-\Delta L}$, as in (27). At this stage, we are able to correctly split the action by writing $\Delta=\Delta_{L}+\Delta_{R}$. In the second line, we ”undo” the holographic renormalization to recover the worldline action for each region separately. Crucially, the dictionary (28) implies that $m\neq m_{L}+m_{R}$ but instead,
@@
To have a well-defined variational principle without fixing the metric or dilaton field on-shell, we obtain the following Israel junction condition for sine-dilaton gravity [^5] Similar junction conditions for general dilaton gravity theories, including JT gravity, was derived in [34].
@@
$I_{L}=\frac{1}{2}\int_{\mathrm{bulk-L}}d^{2}x\sqrt{g}\left[\Phi R+2\sin(\Phi)\right]+\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ \Phi K_{L}-im_{L}\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}+\bar{\mu}\int_{\mathrm{shell}}du\sqrt{h}e^{i\Phi/2}.$ (38)
$I_{R}=\frac{1}{2}\int_{\mathrm{bulk-R}}d^{2}x\sqrt{g}\left[\Phi R+2\sin(\Phi)\right]+\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ \Phi K_{R}-im_{R}\int_{\mathrm{shell}}du\sqrt{h}\leavevmode\nobreak\ e^{-i\Phi/2}-\bar{\mu}\int_{\mathrm{shell}}du\sqrt{h}e^{i\Phi/2}.$ (39)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}=\int_{D}d\bar{\mu}\leavevmode\nobreak\ \mathcal{M}(\bar{\mu})\leavevmode\nobreak\ Z_{L}^{\star}(\tau,\bar{\mu},m_{L})\leavevmode\nobreak\ Z_{R}(\beta-\tau,\bar{\mu},m_{R}),$ (40)

### v2

@@ 1 Introduction
As a first step towards understanding bulk physics, in [14, 15] the chord diagrams were argued to be the right language for the emergence of the bulk geometry. Indeed, [15] constructed a two-sided chord Hilbert space defined on the two-sided ”wormhole” Cauchy slice, with extensions to include matter operators, and [16] further analyzed the symmetry algebra acting on this Hilbert space. Later, a series of works [17, 18, 19, 20] proposed a concrete holographic dual of DSSYK, which is called the ”sine-dilaton gravity”. The sine-dilaton gravity was shown to reproduce the dynamics of DSSYK, also referred to as q-Schwarzian quantum mechanics [17, 21, 22], and its canonical quantization reproduces the two-sided chord Hilbert space of [16] . Crucially in this construction, the length in the bulk becomes positive and discrete after gauging the shift symmetry of the conjugate momentum [19]. Besides the connection to DSSYK, the sine-dilaton gravity itself has some very unusual features for a gravitational theory, such as bounded spectrum and the resulting good UV properties, and therefore merits investigation beyond its connection to the SYK model.
In this work, we take the next step in exploring this duality, by including bulk matter fields and reproducing general correlation functions of boundary-inserted operators from the gravitational perspective. A nontrivial example is the out-of-time ordered correlator (OTOC). In DSSYK, the OTOC has been calculated using the Wick contractions in the chord diagram [13], where it gives the 6j-symbol of the quantum group $\mathcal{U}_{q}(su(1,1))$. Notably, this 6j-symbol leads to sub-maximal chaos, which is related to the ”fake thermal circle” of DSSYK [16]. On the gravity side, a direct calculation of OTOC is more subtle. For example, previous OTOC calculations in JT gravity rely on boundary-bulk mappings or bootstrap techniques [23, 24, 25, 26, 27]. However, in sine-dilaton gravity, a more straightforward bulk calculation can be expected – since the bulk geometry is discrete, the separation of different bulk regions in the OTOC contour is less ambiguous, as we will see below.
@@ 1.1 Summary
The motivation to investigate this procedure originates from a natural bulk perspective: a boundary-to-boundary particle worldline on the Euclidean $\mathrm{AdS}_{2}$ disk divides the disk into two regions and can be considered as an end-of-the-world (EOW) brane for each region. When the time direction is chosen such that the Cauchy slices are orthogonal to the EOW brane, this brane acts as a defect operator, leading to a twisted Hilbert space, which we denote by $\mathcal{H}_{\mathrm{\Delta}}$, with $\Delta$ being one of the brane parameter corresponding to the scaling dimension of the inserted boundary matter operator. The canonical quantization result of this twisted Hilbert space was established in [20], where the resolution of the identity operator takes the form
$\textbf{Twisted Hilbert space {$\mathcal{H}_{\mathrm{\Delta}}$}}:\quad\mathds{1}=\int_{0}^{\pi}d\theta\penalty 10000\ \rho_{\Delta,\bar{\alpha}}(\theta)|\theta\rangle_{\Delta,\bar{\alpha}}{{}_{\Delta,\bar{\alpha}}}\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|.$ (1)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}\quad=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}.$ (2)
@@
    $|\Delta;\theta_{L},\theta_{R}\rangle_{0}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=|{\theta_{L}}\rangle\otimes^{\theta_{L}}|{\theta_{R}}\rangle_{\Delta,\theta_{L}},$ (3)
@@
    $|\Delta;\theta_{L},\theta_{R}\rangle_{1}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\int_{0}^{\pi}d\bar{\alpha}\,\rho(\bar{\alpha})\,|{\theta_{L}}\rangle_{\Delta_{L},\bar{\alpha}}\otimes^{\bar{\alpha}}|\bar{\alpha}\rangle\otimes^{\bar{\alpha}}|\theta_{R}\rangle_{\Delta_{R},\bar{\alpha}}$ (5)
@@
$|\Delta_{1},...,\Delta_{n};\theta_{1},...,\theta_{n+1}\rangle=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$},$ (7)
@@
$\mathrm{OTOC}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}.$ (10)
@@
$\int_{0}^{\pi}\int_{0}^{\pi}d\bar{\alpha}_{1}d\bar{\alpha}_{2}\,\rho(\bar{\alpha}_{1})\,\rho(\bar{\alpha}_{2})\quad\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}=\hbox{$\vbox{\hbox{\resizebox{}{}{{\hbox{\set@color{}}}}}}$}.$ (11)
@@ 1.2 Sine-dilaton gravity setup
$\int\mathcal{D}g\mathcal{D}\Phi\penalty 10000\ \mathrm{exp}\left(\frac{1}{2|\log q|}\left\{\frac{1}{2}\int d^{2}x\sqrt{g}\left(\Phi R+2\sin(\Phi)\right)+\int_{\mathrm{bdy}}du\sqrt{h}\penalty 10000\ \Phi K-i\int_{\mathrm{bdy}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}\right\}\right),$ (12)
@@
$\sqrt{h}\penalty 10000\ e^{i\Phi/2}\Big|_{\Phi=\Phi_{\mathrm{bdy}}}=\frac{i}{2|\log q|},\quad\Phi_{\mathrm{bdy}}=\frac{\pi}{2}+i\infty.$ (13)
@@
$L=\int_{\mathrm{slice}}ds\penalty 10000\ e^{-i\Phi/2}=\int_{\mathrm{slice}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}.$ (17)
@@
$\mathds{1}=\int_{0}^{\pi}d\theta\penalty 10000\ \rho(\theta)|\theta\rangle\langle\theta|=\sum_{n=0}^{\infty}|n\rangle\langle n|,$ (21)
@@
$Z_{\mathrm{SD}}(\beta)$ $=\langle n=0|e^{-\beta\hat{H}_{\mathrm{SD}}}|n=0\rangle=\int_{0}^{\pi}d\theta\penalty 10000\ \rho(\theta)\penalty 10000\ e^{\beta\frac{\cos(\theta)}{\hbar}},$ (25)
@@ 2 Matter fields in sine-dilaton gravity
$\langle\mathcal{O}(x_{1})\mathcal{O}(x_{2})\rangle=\int\mathcal{D}x\exp\left(-\frac{m}{\hbar}\int_{x(u_{i})=x_{1}}^{x(u_{f})=x_{2}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}\,+\mathrm{c.t.}\right)=\ \exp\left(-\Delta L(x_{1},x_{2})\right),$ (27)
where we emphasize that the action $\frac{m}{\hbar}\int_{x(u)}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}$ is formally divergent as the bulk operator approaches the boundary, and $L(x_{1},x_{2})$ is a suitably renormalized geodesic length between the boundary points $x_{1}$ and $x_{2}$ in $\mathrm{AdS}_{2}$, obtained after adding the appropriate counter term (c.t.) [^2] The precise form of the counter term, as well as the procedure for obtaining the effective action on the geodesic, is currently unknown. The difficulty in identifying it arises from the fact that the boundary of sine-dilaton gravity is essentially non-geometric and singular [33], as $\mathrm{AdS}_{2}$ patch of sine-dilaton gravity is defined along a complex contour that extends to $i\infty$ and the classical curvature diverges. In this paper, we take the conjectured dictionary (28), inspired by the quantum group structure and the JT gravity result [34], where the effective action $e^{-\Delta L(x_{1},x_{2})}$ is in fact the exact result for any $\Delta$, as the definition of this counter term. It would be interesting and important to investigate this point in the future. For recent explorations of the dictionary, see [22, 35], and $\Delta$ is the scaling dimension of the operator inserted at the boundary. The relation of the scaling dimension to the particle mass is conjectured to be [20, 17]
@@ 2.1 Splitting a particle worldline in $\mathrm{AdS}_{2}$
$-im\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}+\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ \Phi K_{L}+\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ \Phi K_{R}.$ (31)
@@
$\exp\left(-\frac{m}{\hbar}\int_{x(u)}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}\right)\rightarrow\exp\left(-\Delta L\right)=\exp\left(-(\Delta_{L}+\Delta_{R})L\right)$
$\rightarrow\exp\left(-\frac{m_{L}}{\hbar}\int_{x(u)}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}-\frac{m_{R}}{\hbar}\int_{x(u)}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}\right)$ (32)
where in the first line, the arrow indicates the use of holographic renormalization and summing over all possible shell locations to obtain the finite action $e^{-\Delta L}$, as in (27). At this stage, we are able to correctly split the action by writing $\Delta=\Delta_{L}+\Delta_{R}$. [^5] One can choose any decomposition of $\Delta=\Delta_{L}+\Delta_{R}$, and any observables only depends on $\Delta_{L}+\Delta_{R}$. This provides a non-trivial consistency check of our calculation later. In the second line, we ”undo” the holographic renormalization to recover the worldline action for each region separately. Crucially, the dictionary (28) implies that $m\neq m_{L}+m_{R}$ but instead,
@@
To have a well-defined variational principle without fixing the metric or dilaton field on-shell, we obtain the following Israel junction condition for sine-dilaton gravity [^6] Similar junction conditions for general dilaton gravity theories, including JT gravity, were derived in [37].
@@
$I_{L}=\frac{1}{2}\int_{\mathrm{bulk-L}}d^{2}x\sqrt{g}\left[\Phi R+2\sin(\Phi)\right]+\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ \Phi K_{L}-im_{L}\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}+\bar{\mu}\int_{\mathrm{shell}}du\sqrt{h}e^{i\Phi/2}.$ (38)
$I_{R}=\frac{1}{2}\int_{\mathrm{bulk-R}}d^{2}x\sqrt{g}\left[\Phi R+2\sin(\Phi)\right]+\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ \Phi K_{R}-im_{R}\int_{\mathrm{shell}}du\sqrt{h}\penalty 10000\ e^{-i\Phi/2}-\bar{\mu}\int_{\mathrm{shell}}du\sqrt{h}e^{i\Phi/2}.$ (39)
@@
$\langle\mathcal{O}(\tau)\mathcal{O}(0)\rangle_{\beta}=\int_{D}d\bar{\mu}\penalty 10000\ \mathcal{M}(\bar{\mu})\penalty 10000\ Z_{L}^{\star}(\tau,\bar{\mu},m_{L})\penalty 10000\ Z_{R}(\beta-\tau,\bar{\mu},m_{R}),$ (40)